# MoMentS: A Comprehensive Multimodal Benchmark for Theory of Mind

### v1

@@ 7 Conclusion
We adopt a multiple-choice QA format in MoMentS to streamline annotation and ensure consistent evaluation. While this design supports scalable benchmarking, it limits analysis of lower-level behavioral cues such as turn-taking, speech acts, or gesture dynamics as we do not provide annotations on them. Investigating the relation between these cues and specific ToM abilities remains an important direction for future work. Additionally, MoMentS uses static video data, which does not capture model performance in interactive or dynamic social environments. Extending evaluation to such settings is a promising but currently challenging task, as it would require reliably simulating complex, multimodal human behaviors. Finally, although using multiple annotators per question could reduce subjectivity, resource constraints limited us to one annotator per question. To mitigate this, we incorporated peer-checking during distractor creation and conducted multiple rounds of author review to ensure data quality and consistency.
@@ A.4 Evaluation on ASR quality
We evaluate the ASR pipeline using different base Whisper models on a subset of 50 human-transcribed videos, reporting global Word-Error Rate (WER) and Diarization Error Rate (DER. For global WER we concatenate each file’s reference and ASR transcripts lower-casing and punctuation removal and computing $\text{WER}=(S+I+D)/N$, where $S$, $I$, and $D$ are the numbers of substituted, inserted, and deleted words, and $N$ is the total number of reference words. For DER, we evaluate only within spans where the reference marks speech. The score is $\text{DER}=(T_{\text{missed}}+T_{\text{confusion}})/T_{\text{ref}}$, where $T_{\text{missed}}$ is reference speech with no ASR coverage, $T_{\text{confusion}}$ is overlapped speech attributed to the wrong mapped speaker, and $T_{\text{ref}}$ is the total duration of speech in the reference annotation. We report these in Table 6, while Whisper large-v3 scores the lowest average global WER, in practice we notice that it failed to transcribe some of the videos. This does not happen with large-v2, whose DER is the lowest; because of this, we opted for the latter as the chosen model for transcribing audio for the Video LLMs.
@@
base 0.362 0.412
large-v2 0.206 0.363
large-v3 0.166 0.409
@@ A.5 Ablation on number of frames
As shown in Table 7, increasing the number of frames does not lead to consistent improvements. In several cases, performance actually drops, likely due to redundancy or context saturation. Based on these results, we use 64 frames for all main evaluations in the paper.
$[t_{0},t_{j}]$ $[t_{i},t_{j}]$
Model $T$ $VT$-64 $VT$-96 $T$ $VT$-64 $VT$-96
@@
Table 7: Global accuracy on a subset of 1,500 MoMentS samples using only transcripts ($T$), and transcripts plus 64 or 96 frames ($VT$-64 and $VT$-96). Results are reported for both the Full ($[t_{0},t_{j}]$) and Focused ($[t_{i},t_{j}]$) Context Windows. We mark in bold the highest increase over $T$ between 96 and 64 frames.
A.6 Dataset Statistics and Annotation Cost
The main annotation batch involved 16 participants: 12 undergraduate students in psychology and social sciences, two computer scientists, and two clinical psychologists. 12 of them were female and 4 male, all of them between 20 and 30 years old. Twelve of the annotators were from Canada, and the remaining were from Mexico. All participants were explained the purpose of their annotations in an onboarding session.
@@ A.6 Dataset Statistics and Annotation Cost
ToM Ability # Questions
Emotions 599
Beliefs 379
Desires 386
Intentions 1026
Percepts 316
Knowledge 329
NLC 222
Table 8: Number of questions associated with each ToM ability.
@@
MoMentS contains 2,344 questions across 168 short films, the majority of which are in English (144). We also include a subset of 11 films in other languages. Table 9 reports the number of videos per language.
In Table 2, we report the average question length, average answer lengths, and durations of the full and focused context windows. We also display the distributions of lengths for answers, focused, and full context windows in Figures 6, Figure 7, and Figure 8, respectively.
A.7 Copyright and License
@@ A.8 Ethical Considerations
Figure 6: Distribution of lengths for correct answers and distractors.
Figure 7: Length distribution of the Focused Context Windows.
Figure 8: Length distribution of the Full Context Windows.
A.9 Guidelines for Question and Distractor Annotation

### v2

@@ 7 Conclusion
We adopt a multiple-choice QA format in MoMentS to streamline annotation and ensure consistent evaluation. While this design supports scalable benchmarking, it limits analysis of lower-level behavioral cues such as turn-taking, speech acts, or gesture dynamics as we do not provide fine-grained annotations on them. Investigating the relation between these cues and specific ToM abilities remains an important direction for future work. Additionally, MoMentS uses static video data, which does not capture model performance in interactive or dynamic social environments. Extending evaluation to such settings is a promising but currently challenging task, as it would require reliably simulating complex, multimodal human behaviors. Finally, although using multiple annotators per question could reduce subjectivity, resource constraints limited us to one annotator per question. To mitigate this, we incorporated peer-checking during distractor creation and conducted multiple rounds of author review to ensure data quality and consistency.
@@ A.4 Evaluation on ASR quality
English Non-English
Model $t_{0},t_{j}] [$$t_{i},t_{j}] [$$t_{0},t_{j}] [$$t_{i},t_{j}]$
$T$ VTVT $T$ VTVT $T$ VTVT $T$ VTVT
⚫ LLaVA-Video-7B 47.59 49.11 46.53 51.39 43.90 50.68 40.38 55.28
⚫ InternVL2.5 8B 46.33 46.84 45.06 50.84 44.17 45.53 47.43 56.91
⚫ LongVA-7B-DPO 41.37 43.85 41.11 44.71 38.75 46.34 43.36 43.36
⚫ Qwen2.5 VL 7B 40.71 38.23 37.92 44.15 44.72 37.13 40.92 45.26
⚫ LLaVA-Video-72B 62.99 66.08 62.58 67.59 64.23 65.31 59.35 68.02
⚫ InternVL2.5 78B 53.72 61.11 52.35 60.46 51.22 60.98 51.49 66.94
⚫ Average 48.78 50.87 47.59 53.19 47.83 50.99 47.15 55.96
$T$ VTVT $T$ VTVT $T$ VTVT $T$ VTVT
▲ Qwen2.5-Omni-7B 46.38 52.71 45.57 54.99 49.32 57.18 47.15 62.60
▲ VideoLLaMA2-7B-AV 36.76 40.66 38.58 42.99 42.28 42.55 37.40 43.90
▲ MiniCPM-o 2.6 (8B) 47.70 46.99 49.37 48.51 47.43 54.47
▲ Average 43.61 46.68 43.71 49.11 46.70 49.86 43.99 53.66
AA VAVA AA VAVA AA VAVA AA VAVA
▲ Qwen2.5-Omni-7B 44.30 52.15 48.91 54.28 44.99 55.56 46.07 62.60
▲ VideoLLaMA2-7B-AV 33.77 41.92 33.82 42.99 36.59 44.44 36.31 46.88
▲ MiniCPM-o 2.6 (8B) 39.54 ** 40.00 48.10 39.02 ** 37.67 49.05
▲ Average 39.21 47.04 40.91 48.46 40.20 50.00 40.02 52.85
AA AA AA AA
▼ Kimi-Audio-7B 31.95 48.41 30.08 49.59
▼ Qwen2-Audio-7B 35.54 35.65 29.00 34.42
▼ Average 33.75 42.03 29.54 42.01
Table 6: Global accuracy of different Video LLMs (⚫), Audiovisual LLMs (▲), and Speech LLMs (▼) across English and non-English videos.
@@
We evaluate the ASR pipeline using different base Whisper models on a subset of 50 human-transcribed videos, reporting global Word-Error Rate (WER) and Diarization Error Rate (DER). For global WER we concatenate each file’s reference and ASR transcripts lower-casing and punctuation removal and computing $\text{WER}=(S+I+D)/N$, where $S$, II, and $D$ are the numbers of substituted, inserted, and deleted words, and $N$ is the total number of reference words. For DER, we evaluate only within spans where the reference marks speech. The score is $\text{DER}=(T_{\text{missed}}+T_{\text{confusion}})/T_{\text{ref}}$, where $T_{\text{missed}}$ is reference speech with no ASR coverage, $T_{\text{confusion}}$ is overlapped speech attributed to the wrong mapped speaker, and $T_{\text{ref}}$ is the total duration of speech in the reference annotation. We report these in Table 7, while Whisper large-v3 scores the lowest average global WER, in practice we notice that it failed to transcribe some of the videos. This does not happen with large-v2, whose DER is the lowest; because of this, we opted for the latter as the chosen model for transcribing audio for the Video LLMs.
@@
base 36.2 41.2
large-v2 20.6 36.3
large-v3 16.6 40.9
@@ A.5 Ablation on number of frames
As shown in Table 8, increasing the number of frames does not lead to consistent improvements. In several cases, performance actually drops, likely due to redundancy or context saturation. Based on these results, we use 64 frames for all main Video-LLMs evaluations in the paper.
$t_{0},t_{j}] [$$t_{i},t_{j}]$
Model $T$ VTVT-64 VTVT-96 $T$ VTVT-64 VTVT-96
@@
Table 8: Global accuracy on a subset of 1,500 MoMentS samples using only transcripts ($T$), and transcripts plus 64 or 96 frames (VTVT-64 and VTVT-96). Results are reported for both the Full ($t_{0},t_{j}]$) and Focused ($t_{i},t_{j}]$) Context Windows. We mark in bold the highest increase over $T$ between 96 and 64 frames.
A.6 Performance comparison in English and Non-English videos.
As noted in Appendix A.7, MoMentS includes a subset of non-English videos. Table 6 compares performance on English-only and non-English clips. We find no substantial drops in accuracy for non-English videos; in fact, models with visual inputs often perform better in this setting. A possible explanation for this is that most of the non-English videos include subtitles in the frames, which may support the temporal grounding of dialogues.
A.7 Dataset Statistics and Annotation Cost
The main annotation batch involved 16 participants: 12 undergraduate students in psychology and social sciences, two computer scientists, and two clinical psychologists. 12 of them were female and 4 male, all of them between 20 and 30 years old. Twelve of the annotators were from Canada, and the remaining were from Mexico. All participants were explained the purpose of their annotations in an onboarding session. The annotation process received approval from the MBZUAI Ethics Review Board.
@@ A.7 Dataset Statistics and Annotation Cost
MoMentS contains 2,335 questions across 168 short films, the majority of which are in English (144). We also include a subset of 24 films in other languages. Table 9 reports the number of videos per language.
In Table 2, we report the average question length, average answer lengths, and durations of the full and Focused Context Windows. We also display the distributions of lengths for answers, Focused, and Full context windows in Figures 6(a), Figure 6(b), and Figure 6(c), respectively.
A.8 Copyright and License
@@ A.9 Ethical Considerations
ToM Ability # Questions
Emotions 599
Beliefs 379
Desires 386
Intentions 1026
Percepts 316
Knowledge 329
NLC 222
Table 10: Number of questions associated with each ToM ability.
@@
(a) Distribution of lengths for correct answers and distractors.
(b) Length distribution of the Focused Context Windows.
(c) Length distribution of the Full Context Windows.
Figure 6: Histograms of different the statistics reported in Table 2.
A.10 Guidelines for Question and Distractor Annotation